Home‎ > ‎Crew 134‎ > ‎

126-CommandersReport

Commander Report
01/26/2014
Ashley Dale

Today was a very productive day. Ewan and Kai finally debugged the rover. The rover is fully 
operational! We conducted our first full rover EVA today, and tested out the stereo-cam system. 
Susan confirmed another tele-anaesthesiology / tele-surgery study to be conducted in a few days with the Antarctic base, Concordia, and experts from around the world. The radio signal quality data 
collection is going swiftly. Vibha has the hydrogen-from-soil extraction hardware functioning well. Michaela collected more endolith samples. Sue Ann is collecting great data to narrow down the design envelope for the next generation of haptic glove technologies she’s working on. Kai is in full swing with generating b-roll and editing crew interviews and outreach video packages.

We’ve got a reporter from CNN with us for tomorrow, which should hopefully be just a minor distraction from the work we’re doing. But we are hopeful that what he produces will be of benefit 
for the underlying message MDRS represents, along with the bi-daily videos we have featured on 
space.com

The formation of this crew has been a year in the making. I am proud of what has been achieved 
already. A lot of good will come from this group in the next week, and the coming year.

-----------------------

CommanderCheckIn Report
01/26/2014
Ashley Dale
Crew Physical Health:
Physically everyone is in good condition. Vibha's ankle still plays up 
a little when straining it too much. Susan still has some stiffness in 
places but is far better than yesterday.
   
Eva Depart Time:
10h10
Eva Return Time:
12h30
Narrative Of Field Mission Results:
Radio Signal Quality EVA - success. Walked several kilometers north and west of HAB. 
Eva Depart Time:
10h30
Eva Return Time:
12h15
Narrative Of Field Mission Results:
Success collection of samples.
Eva Depart Time:
15h30
Eva Return Time:
17h45
Narrative Of Field Mission Results:
Successful rover field-testing. Made it to rover site 1 and tested stereo-camera system. 
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:
   The trailer water tank is very low. Tank outside HAB is just about full. Strong smell of propane coming from GreenHAB - more so than normal. Made Ewan/Kai feel ill while they were working on there for several hours.
Rep Sched:
n/a   
Tomorrow Plan:
Reporter from CNN here for the day. Rover field-testing. Further endolith fieldwork. Further work planned with head-mounted EEG scanners.  
Inventory:
   n/a
Support Req:
Trailer water tank is low
Miscelleneous:
n/a